# DReSS: Data-driven Regularized Structured Streamlining for Large Language Models

### v1

@@ 4.2 Performance Comparison
As shown in Table 1, DReSS demonstrates superior performance across both generation and zero-shot tasks achieving lowest perplexity and highest average accuracy across all models. For instance, on LLaMA2-7B, its perplexity is 20% lower than the second-best method, LLM Surgeon. Notably, on OPT-13B, compared to the dense model, DReSS experiences only a 1% drop in average accuracy. The performance difference between the $\ell_{2}$-norm and $\ell_{1}$-norm is tiny. In Figure 3, we present the ratio of the absolute values of model parameters and activation values after and before regularization of the selected portions. The results show that, in most intermediate layers, the parameters and activation values after regularization are reduced to 30% of their original values, which indicates that the amount of information contained in this part of the parameter space has decreased. Overall, regularization may help transfer important information hidden in the pruned portions to the remaining parts.
4.3 Acceleration Comparison
@@
4.4 Robustness to Varying Pruning Ratios.
Keeping all other settings consistent with Section 4.2, we extend the pruning ratio from 20% to 60%. The perplexity of LLaMA2-7B on WikiText-2 under different methods are shown in Figure 4. DReSS significantly outperforms other methods across various pruning ratios. At a pruning ratio of 20%, DReSS performs similarly to the dense model. When the pruning ratio is up to 60%, SLEB collapses, while DReSS maintains relatively low perplexity compared to other pruning methods. This demonstrates that DReSS maintains robust performance even under extreme pruning ratios, enabling structured pruning to achieve high sparsity levels and unlocking significant potential for model acceleration.
@@ 4.5 Dependency on Datasets
Since DReSS relies on data-driven regularization, we investigate its potential dataset dependence. We evaluated perplexity for four methods on WikiText-2, using calibration and RFT data selected from Alpaca, WikiText-2, PTB, and C4. To ensure fairness, we randomly selected 1,000 samples from each dataset, with other settings consistent with Section 4.2. As shown in Figure 5, DReSS consistently outperforms the other methods across datasets, demonstrating its robustness and scalability.
@@ 4.7 Ablation Study
Effectiveness of Regularization As shown in Table 4, both random pruning and full-parameter regularization significantly increase perplexity on WikiText-2 and reduce average accuracy at a 25% pruning ratio for LLaMA2-7B and LLaMA2-13B. This underscores the importance of applying regularization only to the pruned components. Additionally, RFT after pruning has minimal impact on perplexity and average accuracy, indicating that the RFT process is optional.
Table 4: Ablation results on LLaMA2-7B and LLaMA2-13B. ‘FPR’ refers to full parameter regularization, ‘R’ refers regularization. “w/o” denotes variants with specific components removed.
@@
Data Ratio of Regualrization and RFT With all other settings consistent with Section 4.2, we only varied the data ratio, and the results are shown in Table 5. It can be observed that when the ratio is 3:1, the model achieves the lowest perplexity. When the ratio is 4:1, the increase in perplexity is less pronounced, but when the ratio is 1:1, the perplexity increases significantly by 15%. These results underscore the importance of the regularization process relative to the RFT applied after pruning.
Trade Off Between Language Modeling and Regularization Keeping all other settings consistent with Section 4.2, we evaluate the model’s performance by varying $\lambda\in[10^{-5},10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}]$. As shown in Table 6, when $\lambda=10^{-5}$, the regularization loss is too small to effectively shift important values, leading to a significant increase in perplexity after pruning. In contrast, when $\lambda=10^{-1}$, excessive regularization compromises the model’s language modeling ability, causing a notable performance decline post-pruning. Optimal performance is achieved at $\lambda=10^{-3}$, demonstrating the importance of balancing language modeling loss and regularization loss for effective pruning.
@@ 5 Conclusion
In this paper, we introduce a novel pruning paradigm: first applying regularization to the pruned components, followed by pruning and then RFT. Unlike traditional paradigms that first prune based on importance and then apply RFT, DReSS transfers critical information from the pruned parameter space to the remaining components during the regularization, effectively mitigating the irreversible performance degradation caused by information loss. DReSS demonstrates superior performance in both generation and zero-shot tasks, significantly outperforming existing pruning methods. For instance, DReSS surpasses the powerful LLM Surgeon by 20% in perplexity on LLaMA2-7B. On OPT-13B, with a pruning ratio of 25%, the average accuracy drops by only 1%, while achieving a 1.41× speedup compared to the dense model. Moreover, DReSS requires only 25% of the data to achieve comparable performance to previous methods, substantially reducing both data and computational costs while minimizing the reliance on RFT.

### v2

@@ 4.2 Performance Comparison
As shown in Table 1, DReSS demonstrates superior performance across both generation and zero-shot tasks achieving lowest perplexity and highest average accuracy across all models. For instance, on LLaMA2-7B, its perplexity is 20% lower than the second-best method, LLM Surgeon. Notably, on OPT-13B, compared to the dense model, DReSS experiences only a 1% drop in average accuracy. The performance difference between the $\ell_{2}$-norm and $\ell_{1}$-norm is tiny. In Figure 3, we present the ratio of the absolute values of model parameters and activation values after and before regularization of the selected portions. The results show that, in most intermediate layers, the parameters and activation values after regularization are reduced to 30% of their original values, which indicates that the amount of information contained in this part of the parameter space has decreased. We also present the changes occurring in the unregularized part in Appendix D. The results show that this part has increased compared to before, indicating that it now accommodates more important information. Overall, regularization transfers important information hidden in the pruned portions to the remaining parts.
4.3 Acceleration Effectiveness
@@
4.4 Robustness to Varying Pruning Ratios
Keeping all other settings consistent with Section 4.2, we extend the pruning ratio from 20% to 60%. The perplexity of LLaMA2-7B on WikiText-2 under different methods are shown in Figure 4. DReSS significantly outperforms other methods across various pruning ratios. At a pruning ratio of 20%, DReSS performs similarly to the dense model. When the pruning ratio is up to 60%, SLEB collapses, while DReSS maintains relatively low perplexity compared to other pruning methods. This demonstrates that DReSS maintains robust performance even under extreme pruning ratios, enabling structured pruning to achieve high sparsity levels and unlocking significant potential for model acceleration. A more detailed exploration of pruning ratios using DReSS is provided in the Appendix E.
@@ 4.5 Dependency on Datasets
Since DReSS relies on data-driven regularization, we investigate its potential dataset dependence. We evaluated perplexity for four methods on WikiText-2, using calibration and RFT data selected from Alpaca, WikiText-2, PTB, and C4. To ensure fairness, we randomly selected 1,000 samples from each dataset, with other settings consistent with Section 4.2. As shown in Figure 5, DReSS consistently outperforms the other methods across datasets, demonstrating its robustness and scalability. More details are shown in Appendix F.
@@ 4.7 Ablation Study
Effectiveness of Regularization: As shown in Table 4, both random pruning and full parameter regularization significantly increase perplexity on WikiText-2 and reduce average accuracy at a 25% pruning ratio. This highlights the importance of applying regularization exclusively to the pruned components. Additionally, RFT after pruning has minimal impact on model performance, indicating that the RFT process is optional.
Table 4: Ablation results on LLaMA2-7B and LLaMA2-13B. ‘FPR’ refers to full parameter regularization, ‘R’ refers regularization. ‘w/o’ denotes the removal of specific components.
@@
Data Ratio of Regualrization and RFT: With all other settings consistent with Section 4.2, we only varied the data ratio, and the results are shown in Table 5. It can be observed that when the ratio is 3:1, the model achieves the lowest perplexity. When the ratio is 4:1, the increase in perplexity is less pronounced, but when the ratio is 1:1, the perplexity increases significantly by 15%. These results underscore the importance of the regularization process relative to the RFT applied after pruning.
Trade-Off Between Language Modeling and Regularization: Keeping all other settings consistent with Section 4.2, we evaluate the model’s performance by varying $\lambda\in[10^{-5},10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}]$. As shown in Table 6, when $\lambda=10^{-5}$, the regularization loss is too small to effectively shift important values, leading to a significant increase in perplexity after pruning. In contrast, when $\lambda=10^{-1}$, excessive regularization compromises the model’s language modeling ability, causing a notable performance decline post-pruning. Optimal performance is achieved at $\lambda=10^{-3}$, demonstrating the importance of balancing language modeling loss and regularization loss for effective pruning.
@@ 5 Conclusion
In this paper, we introduce a novel pruning paradigm: first applying regularization to the pruned components, followed by pruning and then RFT. Unlike traditional paradigms that first prune based on importance and then apply RFT, DReSS transfers critical information from the pruned parameter space to the remaining components during regularization, effectively mitigating the irreversible performance degradation caused by information loss. DReSS demonstrates superior performance in both generation and zero-shot tasks, significantly outperforming existing pruning methods. For instance, DReSS surpasses the powerful LLM Surgeon by 20% in perplexity on LLaMA2-7B. On OPT-13B, with a pruning ratio of 25%, the average accuracy drops by only 1%, while achieving a 1.41× speedup compared to the dense model. Moreover, DReSS requires only 25% of the data to achieve comparable performance to previous methods, substantially reducing both data and computational costs while minimizing the reliance on RFT.
@@
Appendix A Proofs of the Propositions
A.1 Proof of Proposition 3.1
Let us first assume that $A\in\mathbb{R}^{m\times n}$ and $B\in\mathbb{R}^{n\times k}$. We can express $A$ as $\begin{bmatrix}\mathbf{a}_{1}&\mathbf{a}_{2}&\dots&\mathbf{a}_{n}\end{bmatrix}$, where $\mathbf{a}_{i}\in\mathbb{R}^{m\times 1}$, and $B$ as $\begin{bmatrix}\mathbf{b}_{1}^{T}\\ \mathbf{b}_{2}^{T}\\ \vdots\\ \mathbf{b}_{n}^{T}\end{bmatrix}$, where $\mathbf{b}_{i}\in\mathbb{R}^{k\times 1}$. Then, we obtain $C\in\mathbb{R}^{m\times k}$ as shown in Equation 11. It is important to note that the final summation in Equation 11 refers to the summation of $n$ rank-1 matrices, not a scalar summation.
$\begin{split}C=AB=\begin{bmatrix}\mathbf{a}_{1}&\mathbf{a}_{2}&\dots&\mathbf{a}_{n}\end{bmatrix}\times\begin{bmatrix}\mathbf{b}_{1}^{T}\\ \mathbf{b}_{2}^{T}\\ \vdots\\ \mathbf{b}_{n}^{T}\end{bmatrix}=\sum_{i=1}^{n}\mathbf{a}_{i}\mathbf{b}_{i}^{T}\end{split}$ (11)
According to Equation 11, when some columns of $A$ are zero, the rows of $B$ that correspond to the outer product with the zero columns of $A$ can take any values, as the product of zero and any value is zero. These rows can therefore be set to zero.
Applying the above conclusion to the parameters of a transformer layer, if certain columns of $W_{down}^{i-1}$ in the FFN layer of the $(i-1)$th transformer layer are pruned, then, since $W_{Q}^{i}$, $W_{K}^{i}$, and $W_{V}^{i}$ in the $i$th transformer layer are immediately multiplied by the result of $W_{down}^{i-1}$, the corresponding rows in $W_{Q}^{i}$, $W_{K}^{i}$, and $W_{V}^{i}$ can be set to zero. Additionally, if certain columns of $W_{o}^{i}$ in the $i$th transformer layer are pruned, the $W_{up}^{i}$ in the FFN block will immediately multiply with it, so the corresponding rows in $W_{up}^{i}$ should also be pruned. Thus, we have completed the proof of Proposition 3.1.
A.2 Proof of Proposition 3.2
Step 1: Expressing $\ell_{1}$-norm Using Elements.
The objective function in the unconstrained problem is the $\ell_{1}$-norm of the vector $x$, which is defined as:
$||x||_{1}=\sum_{i=1}^{n}|x_{i}|$ (12)
This function aims to minimize the sum of the absolute values of the components of $x$.
Step 2: Reformulating the Constrained Problem
The constrained optimization problem introduces an auxiliary variable $y$, where for each element $i$:
$x_{i}\geq-y_{i}\quad\text{and}\quad x_{i}\leq y_{i}\quad$ (13)
This implies that $y_{i}\geq|x_{i}|$, meaning each element of $y$ serves as an upper bound for the absolute value of the corresponding element in $x$. Consequently, minimizing $|x|_{1}$ is equivalent to minimizing the sum of the elements in $y$. Thus, the objective function is defined as:
$\mathbf{1}^{T}y$ (14)
Thus, minimizing $\mathbf{1}^{T}y$ is equivalent to minimizing the sum of the absolute values of $x$, which is the $\ell_{1}$-norm of $x$.
This transformation allows the optimization problem to be solved without directly involving the absolute value function, resulting in an equivalent constrained optimization problem that can be addressed via backpropagation. Thus, the proof of the Proposition 3.2 is complete.
Appendix B Detailed Implementation
In this part, we first introduce several hyperparameter settings, with the detailed results shown in Table 7. In our experiments, we employ FP16 precision for all evaluated models, including Phi-2, OPT-2.7B, OPT-6.7B, OPT-13B, LLaMA2-7B, and LLaMA2-13B. For all RFT configurations, we set the LoRA rank $r$ to 32, the scaling factor $\alpha$ to 10, and the sequence length to 2048. All other hyperparameters follow the default settings provided in the Hugging Face PEFT package (Mangrulkar et al., 2022). Due to limited computational resources, we set the batch size to 64. In future work, we will further explore the impact of a broader range of batch sizes on the performance of the pruning method. These choices ensure a balance between model adaptability and computational efficiency, leveraging parameter-efficient fine-tuning while maintaining stable optimization dynamics.
Table 7: Implementation Details
Precision LoRA Rank Scaling Factor Max Sequence Length Batch Size Learning Rate Early Stop Threshold Min Delta FP16 32 10 2048 64 2e-5 5 1e-4
To ensure a fair comparison between DReSS and other methods, we maintain consistency in the data used across all approaches. Specifically, the data used by DReSS for regularization, by LLM Surgeon for periodic updates of model weights and structures, by SliceGPT for selecting channel importance, and by SLEB for identifying crucial transformer layers are identical. Furthermore, we ensure that the data employed during the RFT process is consistent across all methods, thereby enabling a controlled and equitable evaluation framework. Following previous works (Ashkboos et al., 2024; Song et al., 2024), for the comparison unstructured pruning methods like Mangniude, Wanda, and SparseGPT, we ensure that the data used to compute the importance of individual weights is the same as the data used by DReSS for regularization.
Appendix C Results of Comparison with Unstructured Pruning Methods
We compare DReSS at a 25% pruning ratio with unstructured pruning methods employing 2:4 sparsity (where two out of every four consecutive parameters are set to zero) and present the experimental results in Table 8. For unstructured pruning methods, only calibration data is required. To ensure consistency, all methods utilize a total of 1,000 samples, randomly selected from WikiText-2. The models used in these experiments consists of the OPT and LLaMA model families.
Compare to structured pruning methods, the primary advantage of unstructured pruning lies in its flexibility, enabling the removal of specific weights while preserving critical information in the remaining parameters of LLMs. Based on the results in Table 8, it can be seen that DReSS with a pruning ratio of 25% significantly outperforms existing unstructured pruning methods with a sparsity of 2:4 on generative and zero-shot tasks, demonstrating the powerful performance of DReSS. The reason DReSS outperforms unstructured pruning in comparisons lies in the regularization process before pruning. This process transfers important information hidden in the parameter space to the remaining parts of the model, preventing irreversible performance degradation caused by direct pruning and the resulting loss of information.
Table 8: Performance comparison with unstructured pruning methods. ‘PPL’ refers to the perplexity on WikiText-2. The accuracy is reported on five zero-shot tasks: PIQA, WinoGrande, HellaSwag, ARC-e, and ARC-c, as well as the average accuracy. The best result is highlighted in bold, and the second-best in underlined. DReSS-$\ell_{2}$ denotes the use of the $\ell_{2}$-norm, while DReSS-$\ell_{1}$ denotes the use of the $\ell_{1}$-norm.
Model Method Pruning Ratio PPL ($\downarrow$) PIQA WinoGrande HellaSwag ARC-e ARC-c Avg_Acc OPT-2.7B Dense 0% 12.46 74.81 61.01 60.58 54.42 31.14 56.39 Magnitude 2:4 265 57.46 42.19 35.62 32.58 16.42 36.85 Wanda 2:4 16.54 64.32 57.84 41.63 43.47 23.44 46.14 SparseGPT 2:4 13.48 68.58 56.72 45.36 46.89 25.92 48.69 DReSS-$\ell_{2}$ 25% 12.52 69.12 58.98 49.10 50.62 28.15 51.20 DReSS-$\ell_{1}$ 25% 12.48 68.89 58.33 48.62 51.67 27.55 51.01 OPT-6.7B Dense 0% 10.85 76.39 65.19 67.16 60.14 34.64 60.70 Magnitude 2:4 969 62.16 48.74 43.48 41.57 20.63 43.32 Wanda 2:4 13.29 71.76 60.22 54.20 51.18 28.33 53.14 SparseGPT 2:4 11.55 74.21 60.77 57.25 53.03 29.44 54.94 DReSS-$\ell_{2}$ 25% 11.06 74.26 63.12 61.26 56.64 31.77 57.41 DReSS-$\ell_{1}$ 25% 11.74 73.86 64.28 60.45 56.36 32.07 57.40 OPT-13B Dense 0% 10.12 76.82 64.80 69.81 61.87 35.67 61.79 Magnitude 2:4 1.2e4 64.38 52.19 49.46 44.58 21.25 46.37 Wanda 2:4 13.49 72.25 61.72 57.97 53.96 29.69 55.12 SparseGPT 2:4 11.17 74.16 62.43 59.18 56.27 31.74 56.76 DReSS-$\ell_{2}$ 25% 10.38 74.06 64.57 66.82 61.56 35.33 60.47 DReSS-$\ell_{1}$ 25% 10.56 73.65 64.38 66.46 61.15 34.74 60.07 LLaMA2-7B Dense 0% 5.47 79.11 69.06 75.99 74.58 46.25 69.00 Magnitude 2:4 54.59 64.42 60.93 56.72 59.18 27.31 53.71 Wanda 2:4 11.35 70.84 62.27 55.33 57.58 31.91 55.59 SparseGPT 2:4 8.67 72.14 64.96 58.93 60.90 34.22 58.23 DReSS-$\ell_{2}$ 25% 5.86 73.18 66.49 61.73 65.42 40.86 61.54 DReSS-$\ell_{1}$ 25% 5.81 73.42 65.73 61.92 65.26 39.68 61.20 LLaMA2-13B Dense 0% 4.88 80.47 72.22 79.39 77.48 49.23 71.76 Magnitude 2:4 8.33 70.47 62.04 50.65 62.46 31.74 55.47 Wanda 2:4 8.32 73.94 67.01 63.09 64.31 37.80 61.23 SparseGPT 2:4 7.07 75.46 68.51 65.52 66.04 39.76 63.06 DReSS-$\ell_{2}$ 25% 5.12 76.14 71.22 76.51 73.45 46.87 68.84 DReSS-$\ell_{1}$ 25% 5.16 77.24 70.61 76.75 73.84 45.46 68.78
Appendix D Changes in the Parts Without Regularization
As illustrated in Figure 6, the magnitude of the unregularized parameters exhibits an increase after the application of regularization, suggesting a redistribution of model capacity. This phenomenon indicates that during regularization, critical information, initially encoded in the regularized portion of the model, is partially transferred to the unregularized portion. In contrast, Figure 3 shows a reduction in the magnitude of the regularized parameters after regularization, implying that the imposed constraints effectively suppress the corresponding parameter values, enforcing sparsity or compression in that region.
Collectively, these observations suggest that the regularization process facilitates an implicit redistribution of information across different parameter subsets. Specifically, the regularization term promotes a shift of important model characteristics from the constrained (regularized) portion to the unconstrained (unregularized) portion, thereby preserving essential model knowledge despite the imposed sparsity constraints. Based on this insight, we posit that pruning the regularized portion post-regularization could mitigate information loss, as the core knowledge has already been migrated to the unregularized segment. This pruning strategy effectively reduces parameter redundancy while retaining the model’s language modeling capacity, thereby achieving a more compact and efficient representation without compromising performance.
Figure 6: Ratio of the sum of absolute values of unregularized parameters after LLaMA2-7B regularization to the sum of absolute values of the corresponding parameters before regularization. For example, regularization is not applied to the first 75% of rows in $W_{Q}$, and the absolute sum of the rows after regularization is divided by the sum of the corresponding rows before regularization to obtain the Increased Ratios. This is applied similarly to other matrices such as $W_{K}$, $W_{V}$, $W_{o}$, $W_{gate}$, $W_{up}$, $W_{down}$, and for layer activations $X$, the sum of the unregularized columns is compared to the original sum.
Appendix E Performance of DReSS under Different Pruning Ratios and Datasets
E.1 The Perplexity of DReSS under Different Pruning Ratios and Datasets
In Section 4.2, we utilize 1,000 samples randomly selected from the WikiText-2 dataset to guide the regularization process. Subsequently, we evaluate multiple large language models (LLMs) by measuring changes in perplexity across various generative task datasets, including WikiText-2, Alpaca, PTB, and C4, under pruning rates of 10%, 20%, 30%, 40%, 50%, and 60%. The detailed results, presented in Table 9, indicate that DReSS exhibits greater robustness as model scale increases, suggesting that the proposed method effectively mitigates performance degradation in larger architectures. This highlights the scalability of DReSS and its potential to maintain model efficiency under varying levels of sparsity.
Table 9: Perplexity comparison of DReSS with different pruning ratios. We set the pruning ratios to 10%, 20%, 30%, 40%, 50%, and 60%, and test the perplexity of the OPT and LLaMA2 models on the generation task datasets Alpaca, WikiText-2, PTB, and C4. For DReSS, we use the $\ell_{2}$-norm.
Model Pruning Ratio WikiText-2 Alpaca PTB C4 OPT-2.7B Dense 12.46 11.64 17.97 14.32 10% 12.48 11.71 18.16 14.71 20% 12.51 11.93 19.45 15.42 30% 14.49 12.88 23.58 18.93 40% 18.92 15.46 31.30 24.33 50% 24.57 21.37 45.32 32.15 60% 33.83 31.22 58.73 45.56 OPT-6.7B Dense 10.85 10.27 15.77 12.71 10% 10.91 10.45 16.05 13.18 20% 11.02 10.83 17.58 14.45 30% 12.32 11.46 19.65 16.68 40% 14.26 12.71 25.52 20.94 50% 19.63 15.66 33.78 28.22 60% 28.75 21.39 47.29 39.47 OPT-13B Dense 10.12 9.46 14.52 12.06 10% 10.26 9.65 14.88 12.37 20% 10.47 9.97 15.64 13.15 30% 10.99 10.56 18.81 15.63 40% 11.62 11.25 23.07 19.55 50% 16.15 14.68 29.26 26.21 60% 27.63 20.12 38.59 35.58 LLaMA2-7B Dense 5.47 5.25 7.92 7.26 10% 5.54 5.29 8.06 7.34 20% 5.63 5.37 8.29 7.79 30% 7.39 7.32 9.13 8.46 40% 9.62 8.62 12.37 10.56 50% 13.77 12.59 18.85 15.18 60% 24.38 20.25 29.34 27.37 LLaMA2-13B Dense 4.88 4.63 7.16 6.73 10% 4.94 4.69 7.23 6.96 20% 5.08 4.83 7.61 7.53 30% 5.61 5.42 8.33 8.14 40% 6.25 6.14 9.27 9.05 50% 7.59 7.28 11.58 10.88 60% 12.77 11.78 15.16 13.62
E.2 The Accuracy of DReSS under Different Pruning Ratios on Zero-shot Tasks
To systematically evaluate the performance of DReSS on zero-shot tasks under varying pruning rates, we adopt the experimental setup outlined in Section 4.2, where 1,000 samples are randomly selected from the WikiText-2 dataset to guide the regularization process. We assess the accuracy of different model configurations at pruning rates of 10%, 20%, 30%, 40%, 50%, and 60% across a diverse set of benchmark datasets, including PIQA, WinoGrande, HellaSwag, ARC-e, and ARC-c. The results, summarized in Table 10, provide insights into the impact of sparsity on zero-shot generalization. Notably, the analysis reveals that DReSS maintains competitive performance even at higher pruning rates, demonstrating its effectiveness in preserving reasoning and commonsense understanding across different tasks.
Table 10: Accuracy comparison of DReSS with different pruning ratios. We set the pruning ratios to 10%, 20%, 30%, 40%, 50%, and 60%, and test the accuracy of the OPT and LLaMA2 models on the zero-shot task datasets PIQA, WinoGrande, HellaSwag, ARC-e and ARC-c. For DReSS, we use the $\ell_{2}$-norm. ‘Avg_Acc’ represents the average accuracy.
Model Pruning Ratio PIQA WinoGrande HellaSwag ARC-e ARC-c Avg_Acc OPT-2.7B Dense 74.81 61.01 60.58 54.42 31.14 56.39 10% 70.38 60.12 51.79 52.46 29.78 52.91 20% 69.96 59.47 50.45 51.87 28.62 52.07 30% 64.52 58.76 49.25 50.23 27.52 50.06 40% 61.32 56.27 47.39 48.26 25.57 47.76 50% 59.38 53.46 44.63 46.19 21.66 45.06 60% 52.99 48.74 40.03 43.35 16.25 40.27 OPT-6.7B Dense 76.39 65.19 67.16 60.14 34.64 60.70 10% 75.89 64.61 64.69 58.53 33.47 59.44 20% 75.12 64.23 62.56 57.34 32.95 58.44 30% 72.52 62.63 58.27 54.48 29.99 55.58 40% 67.37 58.59 52.12 49.38 26.87 50.87 50% 61.48 55.62 46.46 47.68 22.97 46.84 60% 55.73 51.52 43.38 45.85 18.62 43.02 OPT-13B Dense 76.82 64.80 69.81 61.87 35.67 61.79 10% 75.46 64.69 68.56 61.79 35.58 61.22 20% 74.78 64.61 67.25 61.66 35.43 60.75 30% 72.62 63.26 65.69 59.64 32.57 58.76 40% 68.67 61.49 62.74 55.98 29.26 55.63 50% 62.19 56.46 58.55 52.15 24.53 50.78 60% 57.43 52.72 51.23 47.62 21.05 46.01 LLaMA2-7B Dense 79.11 69.06 75.99 74.58 46.25 69.00 10% 77.38 68.16 71.28 69.26 43.73 65.96 20% 76.42 67.35 68.26 66.73 41.68 64.09 30% 72.29 64.87 58.53 63.48 39.27 59.69 40% 68.66 61.49 55.42 58.61 36.52 56.14 50% 61.32 55.68 51.79 52.35 31.55 50.54 60% 56.45 51.79 48.96 48.98 27.26 46.69 LLaMA2-13B Dense 80.47 72.22 79.39 77.48 49.23 71.76 10% 79.35 72.15 76.42 76.92 48.57 70.68 20% 78.27 71.98 73.20 75.43 47.62 69.30 30% 75.49 69.73 71.54 72.87 45.41 67.00 40% 73.56 64.46 67.75 68.45 41.28 63.10 50% 68.73 59.82 62.48 60.12 36.14 57.46 60% 62.25 56.27 59.93 53.37 29.77 52.32
Appendix F More Details of the Dependency on Calibration Dataset
When using Alpaca, WikiText-2, C4, and PTB as calibration and RFT data, the perplexity of various methods on Alpaca, C4, and PTB is shown as follows:
Figure 7: Comparison of perplexity on Alpaca using different calibration datasets at a pruning ratio of 25% on LLaMA2-7B.
Figure 8: Comparison of perplexity on C4 using different calibration datasets at a pruning ratio of 25% on LLaMA2-7B.
Figure 9: Comparison of perplexity on PTB using different calibration datasets at a pruning ratio of 25% on LLaMA2-7B.